$SPAD/src/input void.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool void.output
)set message test on
)set message auto off
)clear all
--S 1 of 4
a : Integer
--R
--R                                                                   Type: Void
--E 1
)set message void on
--S 2 of 4
b : Fraction Integer
--R
--R
--R   (2)  "()"
--R                                                                   Type: Void
--E 2
)set message void off
--S 3 of 4
3::Void
--R
--R                                                                   Type: Void
--E 3
--S 4 of 4
% :: PositiveInteger
--R
--R
--RDaly Bug
--R   Cannot convert from type Void to PositiveInteger for value
--R   "()"
--R
--E 4
)spool
)lisp (bye)

———- 4

References^{[1] nothing}

5